OpenStudy (anonymous):

Simplify. square root of 7 over fourth root of 7

OpenStudy (michele_laino):

hint: we can write this: \[\Large {\left( {\sqrt[4]{7}} \right)^2} = \sqrt 7 \]

OpenStudy (michele_laino):

so we have: \[\Large \frac{{\sqrt 7 }}{{\sqrt[4]{7}}} = \frac{{{{\left( {\sqrt[4]{7}} \right)}^2}}}{{\sqrt[4]{7}}} = ...?\] please complete